OpenStudy (anonymous):

if a+b+c+d+e+f=1 then the maximum value of ab+bc+cd+de+ef . Given that a,b,c,d,e,f are positive...???

OpenStudy (anonymous):

1

OpenStudy (anonymous):

$$(a+b+c+d+e+f)^2=1\\a^2+b^2+c^2+d^2+e^2+f^2+2(ab+bc+cd+de+ef)=1$$To minimize \(a^2+b^2+c^2+d^2+e^2+f^2\) we want all to be equal (so the 'weight' is distributed evenly) so presume \(a=b=c=d=e=f=\frac16\) and thus we have \(a^2+b^2+c^2+d^2+e^2+f^2=\frac16\):$$2(ab+bc+cd+de+ef)=\frac56\\ab+bc+cd+de+ef=\frac5{12}$$

OpenStudy (anonymous):

I believe this should work...

OpenStudy (zarkon):

are you sure that is correct

OpenStudy (zarkon):

what is \((a+b+c+d+e+f)^2\) again...

OpenStudy (zarkon):

are you sure that problem states "Given that a,b,c,d,e,f are positive" if that is the case then you are trying to maximize a function over a non-compact set. (you are not guaranteed to find an answer)